# Generating synthetic data for neural operators

### v1

@@ Abstract
Numerous developments in the recent literature show the promising potential of deep learning in obtaining numerical solutions to partial differential equations (PDEs) beyond the reach of current numerical solvers. However, data-driven neural operators all suffer from the same problem: the data needed to train a network depends on classical numerical solvers such as finite difference or finite element, among others. In this paper, we propose a new approach to generating synthetic functional training data that does not require solving a PDE numerically. The way we do this is simple: we draw a large number $N$ of independent and identically distributed ‘random functions’ $u_{j}$ from the underlying solution space (e.g., $H_{0}^{1}(\Omega)$) in which we know the solution lies according to classical theory. We then plug each such random candidate solution into the equation and get a corresponding right-hand side function $f_{j}$ for the equation, and consider $(f_{j},u_{j})_{j=1}^{N}$ as supervised training data for learning the underlying inverse problem $f\rightarrow u$. This ‘backwards’ approach to generating training data only requires derivative computations, in contrast to standard ‘forward’ approaches, which require a numerical PDE solver, enabling us to generate a large number of such data points quickly and efficiently. While the idea is simple, we hope that this method will expand the potential for developing neural PDE solvers that do not depend on classical numerical solvers.
@@ 1 Introduction
The use of deep learning for obtaining numerical solutions to PDE problems beyond the reach of classical solvers shows promise to revolutionize science and technology. Deep learning-based methods have overcome many challenges that classical numerical methods suffer from, among which are the curse of dimensionality and grid dependence.
@@
Methods that attempt to solve PDE problems using deep learning can be split into two main classes: those that solve an instance of a PDE problem by directly approximating the solution (e.g. (6), (25), (20), (33), (32), (3)), and those that consider solutions to a family of PDE problems, also known in the literature as parametric PDEs, through operator learning. In the operator learning approach, the goal is to learn an operator that maps the known parameters to the unknown solution (e.g. (16), (2), (12), (10), (21)). In this paper, we focus on the second class, where we seek solutions to a class of PDE problems instead of an instance. Although the approach we describe is general, we focus on the Fourier Neural Operator (FNO) (10), which is a state-of-the-art neural operator learning method at the time that this paper is being written. We stress that our method is independent of the particular neural operator learning architecture and should remain applicable as a synthetic data generation plug-in as the state-of-the-art architecture evolves.
@@
To the best of our knowledge, classical numerical methods, such as finite differences, finite element (27), pseudo-spectral methods, or other variants, have been used in order to obtain data for training purposes in operator learning. In particular, some works have used finite difference schemes (e.g. (16), (2), (13), (26), (19), (23), (21), (17)). In other works, data has been generated by constructing examples that have a closed-form explicit solution or by using schemes such as finite element, pseudo-spectral schemes, fourth-order Runge-Kutta, forward Euler, etc. (e.g. (15), (30), (8), (22), (11), (14), (24), (28), (4), (29), (18), (31)). While these works are a strong proof-of-concept for neural operators, it is critical to move away from using classical numerical solvers to generate training data for neural operator learning if we want to develop neural operators as a general-purpose PDE solver beyond the reach of classical numerical solvers.
@@
Our approach. Our approach is conceptually simple: suppose we want to train a neural network to learn solutions to a parameterized class of PDE problems of the form (1), which is a form where we know in theory that the solution for any value of the parameter belongs to a Sobolev space which has an explicit orthonormal basis of eigenfunctions and associated eigenvalues. We can generate a large number of synthetic training functions $\{u_{a_{j}}^{k}\}_{j,k}$ in the space as random linear combinations of the first $M$ eigenfunctions, scaled by the corresponding eigenvalues. We can efficiently generate corresponding right-hand side functions $f_{a_{j}}^{k}$ by computing derivatives, $-L_{a_{j}}u_{a_{j}}^{k}=f_{a_{j}}^{k}$. We then use training data $(f_{a_{j}}^{k},a_{j},u_{a_{j}}^{k})_{j=1}^{N}$ to train a neural operator to learn the class of PDEs.
@@
Recall the standard supervised learning setting where the training data are input-output pairs $(x_{j},y_{j}),$ where the input vectors $x_{j}$ are independent and identical draws from an underlying distribution ${\cal D}$, and $y_{j}={\cal G}(x_{j})$, and the goal is to derive an approximation ${\tilde{{\cal G}}}$ with minimal test error $\mathbb{E}_{x\sim{\cal D}}|\tilde{{\cal G}}(x)-{\cal G}(x)|$. In our setting, the function to learn is the operator ${\cal G}:(a,f)\rightarrow u$. Our method of generating $(a_{j},f_{a_{j}}^{k})$ and our approach can be viewed as a best attempt within the operator learning framework to replicate training data within the classical supervised learning setting.
@@
Organization of the paper.The order of this paper goes as follows: in Section 2 we introduce the main idea in more details, in Section 3 we discuss how to determine a space for the unknown functions depending on the problem; and in Section 4 we present some numerical experiments when using our data into known network architectures such as the Fourier Neural Operator (10) (FNO). The types of PDE problems we consider are elliptic linear and semi-linear second order equations with Dirichlet and Neumann boundary conditions, starting with the Poisson equation as a first example and then considering more complicated equations. At the end of this paper, we include an appendix section that includes a description of the mathematical symbols used in this paper. Our data generation code can be found on GitHub under the name synthetic-data-for-neural-operators.
@@ 2 Set-up and Main Approach
where $L=L_{a}$ denotes a differential operator parameterized by $a\in\mathcal{A}$, $\Omega\subset\mathbb{R}^{n}$ is a given bounded domain, and $B(u)$ denotes a given boundary condition. The goal is to find a solution $u$ that solves (1) given $L_{a}$ and $f$. So in a general setting, we wish to learn an operator of the form
@@
where $\mathcal{A},\mathcal{F}$ and $\mathcal{U}$ are function spaces that depend on the specific PDE problem. For example, if we take $Lu=\Delta u$, and $B(u)=u$ then (1) becomes the Poisson equation with zero Dirichlet boundary condition. In that case, we can take $\mathcal{F}=L^{2}(\Omega)$ and $\mathcal{U}=H^{1}_{0}(\Omega)$ and the operator we wish to learn is of the form
@@
The main innovation of our work is in determining the appropriate class of functions for the unknown function $u$ in (1). While from the PDE theory we know that $u$ lives in some Sobolev space (see e.g (7)) in the case of elliptic PDEs, such spaces are infinite-dimensional, and it is unclear at first how to generate functions that serve as good representatives of the full infinite-dimensional space. We propose to generate functions as random linear combinations of basis functions of the corresponding Sobolev space. In the case where we know from theory that the underlying Sobolev space is $H^{1}_{0}(\Omega)$ or $H^{1}(\Omega)$, then we can obtain explicit basis elements that are given by the eigenfunctions of the Laplace operator with Dirichlet and Neumann boundary conditions, respectively.
@@ 3 Drawing synthetic representative functions from a Sobolev space
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded and smooth domain, then the eigenvalue problem for the Laplace operator with zero Dirichlet or Neumann boundary value condition is given by
@@
where $B$ denotes the boundary condition. Here, we consider the two cases of Dirichlet and Neumann boundary conditions, which are given by $B(u)=u$ and $B(u)=\nabla u\cdot\nu$, respectively, where $\nu$ denotes the exterior normal derivative of $\partial\Omega$. From theory, we have that in each of the two cases for $B$, the problem (2) has countably many (nonnegative) eigenvalues, and the corresponding eigenfunctions form an orthonormal basis for $L^{2}(\Omega)$. We reference chapter 8 of (1) on the details of the spectral analysis of the Laplacian.
Let $\{\varphi_{n}\}_{n=1}^{\infty}$ denote the eigenfunctions of problem (2), and let $\{\lambda_{n}\}_{n=1}^{\infty}$ denote the corresponding eigenvalues for the Dirichlet case, $B(u)=u$. Let $\{\psi_{n}\}_{n=1}^{\infty}$ be the eigenfunctions of the problem (2) and $\{\mu_{n}\}_{n=1}^{\infty}$ the corresponding eigenvalues for the Neumann case, $B(u)=\nabla u\cdot\nu$. For the Dirichlet case, the normalized family of the eigenfunctions $\varphi_{n}\!/\sqrt{\lambda_{n}}$ forms an orthonormal basis for $H^{1}_{0}(\Omega)$ equipped with the scalar product $\langle u,v\rangle=\int_{\Omega}\nabla u\cdot\nabla v$. For the Neumann case, if we define $V=\{u\in H^{1}(\Omega):\int_{\Omega}u=0\}$ (which is also a Hilbert space with the same scalar product as $H^{1}(\Omega))$, then the normalized family of the eigenfunctions $\psi_{n}\!/\sqrt{\mu_{n}}$ forms an orthonormal basis for $V$.
3.1 Representative functions in square domains
For the square domain $\Omega=(0,1)^{2}$, sequences of eigenvalues and eigenfunctions of the problem (2) are known explicitly. Namely, for the Dirichlet case, eigenfunctions and the corresponding eigenvalues are given by
$\varphi_{ij}=\sin(i\pi x)\sin(j\pi y),\quad\lambda_{ij}=(i\pi)^{2}+(j\pi)^{2}$ (3)
and in the Neumann case, they are given by
$\psi_{ij}=\cos(i\pi x)\cos(j\pi y),\quad\mu_{ij}=(i\pi)^{2}+(j\pi)^{2}$ (4)
Further normalizing the eigenfunctions, we set
$u_{ij}:=\frac{\sin(i\pi x)\sin(j\pi y)}{\sqrt{\lambda_{ij}}},\quad v_{ij}:=\frac{\cos(i\pi x)\cos(j\pi y)}{\sqrt{\mu_{ij}}}$ (5)
which serve as orthonormal basis elements for $H_{0}^{1}(\Omega)$ and $V$, respectively.
Finally, we now generate our function $u$ by considering linear combinations of the above basis functions in (5) (that is, we consider a truncation up to some positive number $M$ instead of considering the whole series). We use the $u_{ij}$’s when considering elliptic problems with zero Dirichlet boundary condition, and we use the $v_{ij}$ for with zero Neumann condition. Namely, let $a_{ij}\sim N(0,1\!/(i+j))$ and and $M\in\mathbb{N}$, and set
$u(x,y)=\sum_{i,j=1}^{M}a_{ij}u_{ij},\quad v(x,y)=\sum_{i,j=1}^{M}a_{ij}v_{ij}$ (6)
Notice that by construction, functions of the form (6) satisfy the zero Dirichlet and zero Neumann boundary conditions, respectively. In figure 1 we show representative functions of the form (6) when $M=5$ and $M=20$ and notice that for larger $M$, we have larger oscillations as opposed to smaller $M$.
Figure 1: Example plots of functions of the form (6) for $M=5$ and $M=20$.
@@ 3.1 Representative functions in square domains
The above discussion pertains to the case when $\Omega$ is a square domain in $\mathbb{R}^{2}$. Eigenvalues and eigenfunctions of the Laplacian have been studied (e.g. (9)) for domains that are non-square and/or in higher dimensions, suggesting that our approach can be extended to more general domains $\Omega$. In this paper, we restrict to experiments for the case where $\Omega$ is the unit square in $\mathbb{R}^{2}$ for simplicity.

### v3

@@ Abstract
Recent advances in the literature show promising potential of deep learning methods, particularly neural operators, in obtaining numerical solutions to partial differential equations (PDEs) beyond the reach of current numerical solvers. However, existing data-driven approaches often rely on training data produced by numerical PDE solvers (e.g., finite difference or finite element methods). We introduce a "backward" data generation method that avoids solving the PDE numerically: by randomly sampling candidate solutions $u_{j}$ from the appropriate solution space (e.g., $H_{0}^{1}(\Omega)$), we compute the corresponding right-hand side $f_{j}$ directly from the equation by differentiation. This produces training pairs ${(f_{j},u_{j})}$ by computing derivatives rather than solving a PDE numerically for each data point, enabling fast, large-scale data generation consisting of exact solutions. Experiments indicate that models trained on this synthetic data generalize well when tested on data produced by standard solvers. While the idea is simple, we hope this method will expand the potential of neural PDE solvers that do not rely on classical numerical solvers to generate their data.
@@ 1 Introduction
The use of deep learning to obtain numerical solutions to PDE problems beyond the reach of classical solvers shows promise in revolutionizing science and technology. Deep learning-based methods have overcome many challenges that classical numerical methods suffer from, among which are the curse of dimensionality and grid dependence.
@@
Methods that attempt to solve PDE problems using deep learning can be split into two main classes: those that solve an instance of a PDE problem by directly approximating the solution (e.g. [6], [29], [23], [38], [37], [3]), and those that consider solutions to a family of PDE problems, also known in the literature as parametric PDEs, through operator learning. In the operator learning approach, the goal is to approximate the solution operator that maps input functions to the unknown solution (e.g. [18], [2], [14], [11], [24], [22]). In this paper, we focus on the second class, where we seek solutions to a class of PDE problems instead of an instance. Although the approach we describe is general, we focus on the Fourier Neural Operator (FNO) [11], which is a state-of-the-art neural operator learning method at the time that this paper is being written. We stress that our method is independent of the particular neural operator learning architecture and should remain applicable as a synthetic data generation plug-in as the state-of-the-art architecture evolves.
@@
To the best of our knowledge, classical numerical methods, such as finite differences, finite element [32], pseudo-spectral methods, or other variants, have been used to obtain data for training purposes in operator learning. In particular, some works have used finite difference schemes (e.g. [18], [2], [15], [31], [21], [27], [24], [19]). In other works, data has been generated by constructing examples that have a closed-form explicit solution or by using schemes such as finite element, pseudo-spectral schemes, fourth-order Runge-Kutta, forward Euler, etc. (e.g. [17], [35], [8], [26], [13], [16], [28], [33], [4], [34], [20], [36]). While these works are a strong proof-of-concept for neural operators, it is critical to not rely completely on using classical numerical solvers to generate training data for neural operator learning if we want to develop neural operators as general-purpose PDE solvers beyond the reach of classical numerical solvers.
@@
Our approach. Our approach is conceptually simple: suppose we want to train a neural network to learn solutions to a parameterized class of PDE problems of the form (1). If we know that the solution for any value of the parameter belongs to a Sobolev space which has an explicit orthonormal basis of eigenfunctions and associated eigenvalues, we can generate a large number of synthetic training functions $\{u_{a_{j}}^{k}\}_{j,k}$ in the space as random linear combinations of the first so many eigenfunctions, scaled by the corresponding eigenvalues (see Section 3 for more details). We can efficiently generate corresponding right-hand side functions $f_{a_{j}}^{k}$ by computing derivatives, $-L_{a_{j}}u_{a_{j}}^{k}=f_{a_{j}}^{k}$. We then use training data $(f_{a_{j}}^{k},a_{j},u_{a_{j}}^{k})_{j=1}^{N}$ to train a neural operator to learn the class of PDEs.
The general concept of first generating the “unknown” function and then substituting it into an equation is not new; a well-known earlier instance appears in the method of manufactured solutions [25], which is widely used for code verification when developing numerical solvers. By constructing an exact solution, one can compare a numerical approximation against this exact solution. The novelty of our work is that we explore the use of this general concept in a completely different setting: for generating training data for neural PDE solvers, in order to obtain solutions to a family of PDE problems without ever solving the PDE. We do this by coupling manufactured solutions with classical PDE theory to randomly draw unknown functions from the solution space, creating a training dataset that generalizes effectively for operator learning. In particular, our experiments demonstrate that models trained solely on data produced by our “backwards” method still achieve strong performance when tested on data generated by conventional numerical solvers, highlighting generalizations offered by our approach. For more details see Section 4.
@@
Recall the standard supervised learning setting where the training data are input-output pairs $(x_{j},y_{j}),$ where the input vectors $x_{j}$ are independent and identical draws from an underlying distribution ${\cal D}$, and $y_{j}={\cal G}(x_{j})$, and the goal is to derive an approximation ${\tilde{{\cal G}}}$ with minimal test error $\mathbb{E}_{x\sim{\cal D}}|\tilde{{\cal G}}(x)-{\cal G}(x)|$. In our setting, the function to learn is the operator ${\cal G}:(a,f)\rightarrow u$. Our method of generating $(a_{j},f_{a_{j}}^{k})$ and our overall approach can be viewed as a best attempt within the operator learning framework to replicate training data within the classical supervised learning setting.
@@
Organization of the paper. This paper is organized as follows: in Section 2 we introduce the main idea in more detail, in Section 3 we discuss how to determine a space for the unknown functions depending on the problem, and in Section 4 we present numerical experiments using our data in a known network architecture such as the Fourier Neural Operator [11] (FNO). The types of PDE problems we consider are elliptic linear and semi-linear second-order equations with Dirichlet and Neumann boundary conditions, starting with the Poisson equation as a first example and then considering more complicated equations. In Subsection 4.3, we present experiments comparing our method to a more classical approach, where the right-hand side is first generated and the corresponding problem is solved numerically to obtain input-output pairs for training. At the end of this paper, we include an appendix section with a description of the mathematical symbols used in this paper. Our data generation code can be found on GitHub under the repository name synthetic-data-for-neural-operators.
@@ 2 Set-up and Main Approach
where $L=L_{a}$ denotes a differential operator parameterized by $a\in\mathcal{A}$, where $a$ here denotes some abstract parametrization which depends on the type of PDE, see for example subsection 4.4. $\Omega\subset\mathbb{R}^{n}$ is a given bounded domain, and $B(u)$ denotes a given boundary condition. The goal is to find a solution $u$ that solves (1) given $L_{a}$ and $f$. So in a general setting, we wish to learn an operator of the form
@@
where $\mathcal{A},\mathcal{F}$ and $\mathcal{U}$ are function spaces that depend on the specifics of the PDE problem.
For example, if we take $Lu=\Delta u$, and $B(u)=u$ then (1) becomes the Poisson equation with zero Dirichlet boundary condition. In this case, we can take $\mathcal{F}=L^{2}(\Omega)$ and $\mathcal{U}=H^{1}_{0}(\Omega)$ and the operator we wish to learn is of the form
@@
The main innovation of our work is in determining the appropriate class of functions for the unknown function $u$ in (1). While from the PDE theory we know that $u$ lives in some Sobolev space (see e.g [7]) in the case of elliptic PDEs, such space is infinite-dimensional and it is unclear at first how to generate functions that serve as good representatives of the full infinite-dimensional space. We propose to generate functions as random linear combinations of basis functions of the corresponding Sobolev space. In the case where we know from theory that the underlying Sobolev space is $H^{1}_{0}(\Omega)$ or $H^{1}(\Omega)$, then we can obtain explicit basis elements that can be obtained by the eigenfunctions of the Laplace operator with Dirichlet and Neumann boundary conditions, respectively.
@@ 3 Drawing synthetic representative functions from a Sobolev space
In this section, we discuss how to draw representative functions from the solution space in the case of elliptic problems so that they generalize well when used in numerical experiments. See the appendix for the definitions of the function spaces used in this section.
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded open set. Consider the following eigenvalue problem
@@
which is called the Laplace-Dirichlet operator when $B(u)=u$. We say that $\lambda\in\mathbb{R}$ is an eigenvalue to the Laplace-Dirichlet operator if there exists $u\in H_{0}^{1}(\Omega)$ with $u\neq 0$ such that
[…]
If such $u\neq 0$, we say that it is an eigenfunction associated to the eigenvalue $\lambda$. The following theorem is well known in the analysis of PDEs and spectral theory (see Chapter 8 of [1]).
Theorem 1.
The Laplace-Dirichlet operator has countably many eigenvalues $0<\lambda_{1}\leq\lambda_{2}\leq\cdots\lambda_{N}\leq\cdots$. There exists an orthonormal basis $(e_{i})_{i=0}^{\infty}$ of $L^{2}(\Omega)$ such that $e_{i}$ is an eigenfunction of the Laplace-Dirichlet operator, i.e of problem (2), corresponding to the eigenvalue $\lambda_{i}$ for each $i\in\mathbb{N}$. Moreover, $(e_{i}\!/\sqrt{\lambda_{i}})_{i=0}^{\infty}$ is an orthonormal basis of $H_{0}^{1}(\Omega)$ equipped with the scalar product $\langle u,\varphi\rangle=\int_{\Omega}\nabla u\cdot\nabla\varphi$.
This theory extends to more general Hilbert spaces, including different elliptic linear operators or different types of boundary value conditions such as Neumann or mixed (e.g. see Theorem 6.6.1 in [1]).
[…]
3.1 Representative functions in rectangular domains
Eigenfunctions of the Laplace operator are known for the Dirichlet, Neumann, and Robin boundary conditions on rectangular domains of the form $(a_{1},b_{1})\times(a_{2},b_{2})\times\cdots\times(a_{n},b_{n})\subset\mathbb{R}^{n}$. They are also known for some non-rectangular domains, see Figure 9 for an example on a triangular domain. To keep the presentation simple, we will mainly consider Dirichlet ($B(u):=0$) and Neumann ($B(u):=\nabla u\cdot\nu$) boundary conditions on $\Omega:=(0,1)^{2}$.
For the Dirichlet case, the eigenfunctions $e_{ij}$ corresponding to the eigenvalues $\lambda_{ij}$ of problem (2) are given by
$e_{ij}(x,y)=\sin(i\pi x)\sin(j\pi y),\quad\lambda_{ij}=(i\pi)^{2}+(j\pi)^{2},\quad(x,y)\in(0,1)^{2},i,j\in\mathbb{N}.$ (3)
For the Neumann case, they are given by
$e_{ij}(x,y)=\cos(i\pi x)\cos(j\pi y),\quad\lambda_{ij}=(i\pi)^{2}+(j\pi)^{2},\quad(x,y)\in(0,1)^{2},i,j\in\mathbb{N}.$ (4)
Further, normalizing appropriately, we define the following basis elements for $H_{0}^{1}(\Omega)$ and $V$, respectively
$u_{ij}(x,y):=\frac{\sin(i\pi x)\sin(j\pi y)}{\sqrt{(i\pi)^{2}+(j\pi)^{2}}},\quad v_{ij}(x,y)=\frac{\cos(i\pi x)\cos(j\pi y)}{\sqrt{(i\pi)^{2}+(j\pi)^{2}}}$ (5)
@@ 3.1 Representative functions in rectangular domains
Notice that since the $u_{ij}$ are basis elements of $H_{0}^{1}(\Omega)$, for any $w\in H_{0}^{1}(\Omega)$, there exist coefficients $c_{ij}$ such that $w$ can be written precisely as
$w(x,y)=\sum_{i,j=1}^{\infty}c_{ij}u_{ij}(x,y)$
where $c_{ij}=(w,u_{i,j})_{H_{0}^{1}}$.
Keeping the above remark in mind, we generate the unknown functions $u$ (which we assume are from $H_{0}^{1}(\Omega)$ or $V$) as truncated sums of random linear combinations basis functions with prescribed decay in the coefficients. More precisely, let $M,K$ denote positive truncation numbers and let $a_{ij},b_{ij}\sim N(0,1\!/\sqrt{i^{2}+j^{2}})$ generate $u\in H_{0}^{1}(\Omega)$ and $v\in V$ as follows
$u(x,y)=\sum_{i=1}^{M}\sum_{j=1}^{K}a_{ij}u_{ij}(x,y),\quad v(x,y)=\sum_{i=1}^{M}\sum_{j=1}^{K}b_{ij}v_{ij}(x,y)$ (6)
Notice that by construction, functions of the form (6) satisfy zero Dirichlet and zero Neumann boundary conditions, respectively. In experiments, we draw $M$ and $K$ randomly in $\{1,2,\cdots,20\}$, that is to say we use up to the first $20$ basis functions. While these spaces are infinite-dimensional and thus require an infinite number of basis functions, we observe that truncating to the first $M=20$ is sufficient to achieve good generalizations to unseen non-trigonometric $f$ functions.
Remark 2.
While in the eigenvalue problem (2) we are seeking eigenfunctions and eigenvalues of the Laplace-Dirichlet operator, as we can see in experiments later (see e.g. Section 4.2), these functions can be used for nonlinear elliptic PDEs as well. This is because Theorem 1 asserts that $(e_{i}\!/\sqrt{\lambda_{i}})_{i=0}^{\infty}$ is an orthonormal basis for $H_{0}^{1}(\Omega)$, which means that as long as we know apriori that a solution belongs to $H_{0}^{1}(\Omega)$, the same eigenfunctions are ‘good’ representative.
3.2 Representative functions in non-rectangular domains
Our experiments mainly focus on square domains; however, the eigenfunctions and eigenvalues of the Laplacian are also known explicitly for certain specific non-rectangular domains. In particular, the Laplacian eigenfunctions are known for disks, circular annuli, spheres and spherical shells which can generally be described as $\Omega=\{x\in\mathbb{R}^{n}:r<|x|<R\}$, with $n=2,3$, as well as for ellipses and elliptical annuli. In addition, they are known for equilateral triangles, that is when $\Omega=\{(x,y)\in\mathbb{R}^{2}:0<x<1,0<y<\sqrt{3}x,y<\sqrt{3}(1-x)\}$. For more details on eigenfunctions of the Laplacian, see [9].
As for domains that are not of the above type, there could be ways to obtain the eigenfunctions of the Laplacian numerically; however, in this case, we cannot easily take derivatives symbolically – the main reason our method is computationally efficient. A potential way to generalize to any domain shape could be by passing the boundary values as an input during the training phase and asking for the right boundary condition after training is finished to get a prediction; this is an interesting future direction to explore.